OpenStudy (anonymous):

Given p = log q r, q = log r p and r = log p q. Show that pqr = 1

OpenStudy (ajprincess):

p=log q r=>q^p=r q=log r p=>r^q=p r=log p q=>p^r=q q^p=r \((p^r)^p=r\) \(p^{rp}=r\) \(\large (r^q)^{rp}=r\) \(\large r^{qrp}=r=r^1\) Does that help? @Biha

OpenStudy (anonymous):

thank you for the solution.. I been trying so hard to solve this question. Thanks again!

OpenStudy (ajprincess):

welcome:) do u get it?

OpenStudy (anonymous):

Yep, I do :D